OpenStudy (anonymous):

Solve negative two-thirds times a plus two-fifths equals three-fourths for a.

OpenStudy (anonymous):

Oh crap i screwed up my answer

OpenStudy (anonymous):

Do you mean (-2/3 * a) + 2/5 = 3/4 ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Solve negative two-thirds times a plus two-fifths equals three-fourths for a. a equals negative seven-fortieths a equals negative seven-tenths a equals negative twenty-one fortieths a equals negative twenty-one tenths

OpenStudy (anonymous):

thats a b c and d

OpenStudy (anonymous):

Ok first i'd minus 2/5 from both sides, this will give you -2/3 * a = 3/4 - 2/5 therefore -2/3 * a = 0.35 then divide both sides by -2/3, this leaves just a a = 0.35 / (-2/3) a = -0.525 = negative 21/40

OpenStudy (anonymous):

If you don't have calculator to add or subtract fractions you times the denominator of one against the denominator and numerator of the other, do this for both, it will give you a common denominator so you can add/subtract them

OpenStudy (texaschic101):

Did you know that if you multiply everything by the common denominator, it will get rid of the fractions and make them all whole numbers

OpenStudy (anonymous):

^ that's a better way to do it :P